Two cable carry line charge densities: (cables can
Answer #1
Given
linear charge densities of cable A is λA = -3*8.85*10-12 C / m
linear charge density of cable B is    λB = 3*8.85*10-12 C / m
distance between two cables   d = 12 m
electric field generated by cable is
                   E = |λ| / 2πεo R   
    1) electric field generated by cable A at a middle points between the cables is ( r = 6 m )
                    E1 = (-3*8.85*10-12 C / m ) / 2πεo (6m)
                           =    0.0795 N /C       
    2) electric field generated by cable  B at a middle points between the cables is ( r = 6 m )  
                      E2 = (3*8.85*10-12 C / m ) / 2πεo (6m)
                           =     0.0795 N /C   
    total electric field at a middle point between the cables
                  E = E1 + E2
                      = 2*0.0795
                       = 0.159 N /C